OpenStudy (anonymous):

can some help me with factoring quadratics 2y^2+20=14y

OpenStudy (anonymous):

2y^2-14y+20=0 2(y^2-7y+10)=0 2(y-5)(y-2)=0

OpenStudy (saifoo.khan):

Power ranger's correct.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you so much i really dont understand these at all